# Moments of characteristic polynomials and their derivatives for $SO(2N)$ and $USp(2N)$ and their application to one-level density in families of elliptic curve $L$-functions

### v1

@@ 6 A mixed moment of elliptic curve $L$-functions
In [14], Conrey, Farmer and Zirnbauer derived the following ratios conjecture. It is not stated explicitly in that paper, but is mentioned after their equation (6.31) as following from their Conjecture 5.3 and their Lemma 6.8 in analogy with (6.31). There are possibly some typos in that paper, so another resource is [32] where the conjecture is stated at Conjecture 5.1 using a slightly different set notation.
@@
×$\times\oint\left(\frac{M|d|^{2}}{4\pi^{2}}\right)^{\tfrac{1}{2}\sum_{k=1}^{K}\left(w_{k}\right)}\left(\frac{H_{\alpha,\gamma}(w_{1},\dots,w_{K})\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\displaystyle\prod_{k=1}^{K}w_{k}}{\displaystyle\prod_{j=1}^{K}\prod_{k=1}^{K}(w_{j}-\alpha_{k})(w_{j}+\alpha_{k})}\right)\mathrm{d}w_{1}\dots\mathrm{d}w_{K},$ (163)
@@
$Y_{E}(\alpha;\gamma):=$ $\frac{\displaystyle\prod_{j<k\leq K}\zeta(1+\alpha_{j}+\alpha_{k})\zeta(1+2\gamma)}{\displaystyle\prod_{k=1}^{K}\zeta(1+\alpha_{k}+\gamma)}.$ (165)
@@
$A_{E}(\alpha,\gamma)$ ::=:= $\frac{\displaystyle\prod_{1\leq j<k\leq K}(1-11^{-1-\alpha_{j}-\alpha_{k}})(1-11^{-1-2\gamma})}{\displaystyle\prod_{k=1}^{K}(1-11^{-1-\alpha_{k}-\gamma})}$ (168)
×$\times\left(\frac{(1-\lambda(11)11^{-1/2-\gamma})}{\displaystyle\prod_{k=1}^{K}(1-\lambda(11)11^{-1/2-\alpha_{k}})}\right)$
×$\times\displaystyle\prod_{p\neq 11}\left[\frac{\displaystyle\prod_{1\leq j<k\leq K}(1-p^{-1-\alpha_{j}-\alpha_{k}})(1-p^{-1-2\gamma})}{\displaystyle\prod_{k=1}^{K}(1-p^{-1-\alpha_{k}-\gamma})}\right.$
×$\left.\times\frac{1}{1+\frac{1}{p}}\left(\frac{1}{2}\frac{(1-\lambda(p)p^{-1/2-\gamma}+p^{-1-2\gamma})}{\displaystyle\prod_{k=1}^{K}(1-\lambda(p)p^{-1/2-\alpha_{k}}+p^{-1-2\alpha_{k}})}\right.\right.$
@@
×$\times\oint\exp\left(\mathcal{N}_{d}\sum_{k=1}^{K}w_{k}\right)\left(\frac{H_{\alpha,\gamma}(w_{1},\dots,w_{K})\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\displaystyle\prod_{n=1}^{K}w_{n}^{3-2K}}{\prod_{n=1}^{K}(w_{n}-\alpha)(w_{n}+\alpha)}\right)\mathrm{d}w_{1}\dots\mathrm{d}w_{K}.$
This has exactly the same polar structure as (2) - it behaves like the equivalent integral for $SO(2N)$. In the above we have defined $\mathcal{N}_{d}=\log\left(\frac{\sqrt{M}d}{2\pi}\right)$ to make the comparison easier. The $\mathcal{N}$ associated with the largest $d$ in the sum, $\mathcal{N}_{X}$, can be thought to be equivalent to the matrix size $N$ in the random matrix version. Note that in the random matrix case we found both the leading order term and the next to leading order term in $N$, while here, for simplicity, we shall just work with the leading order term in $\mathcal{N}_{X}$. Acting in analogy to the random matrix case at (2) we know that we can shrink the contours of integration onto the poles, with narrow necks connecting them. As we showed in Section 2.1, the integral will cancel on either side of the necks, leaving a sum over three circular contours for each $w$ variable. The only terms which contribute to this sum will have the residue from one $w_{i}$ evaluated at ±$\pm\alpha$, and the rest evaluated at 0. Using this fact, and in analogy with (2.1.3), we get
@@
×$\times\left(\frac{\prod_{1\leq j<k\leq{K-1}}\zeta(1+w_{j}+w_{k})\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}}{\prod_{k=1}^{K-1}\zeta(1+w_{k}+\gamma)\prod_{n=1}^{K-1}(w_{n}-\alpha)(w_{n}+\alpha)}\right)\prod_{k=1}^{K-1}g\left(-\frac{w_{k}}{2}\right)$
×±$\times\oint_{\pm\alpha}A_{E}(w_{1},\ldots,w_{K};\gamma)g\left(\frac{\alpha-w_{K}}{2}\right)\prod_{j=1}^{K-1}\left[\zeta(1+w_{j}+w_{K})(w_{K}^{2}-w_{j}^{2})^{2}\right]$
×$\qquad\times\frac{\exp(\mathcal{N}_{d}w_{K})w_{K}^{3-2K}}{\zeta(1+w_{K}+\gamma)(w_{K}-\alpha)(w_{K}+\alpha)}dw_{K}\;dw_{1}\cdots dw_{K-1}.$
@@
×$\times\left(\frac{\prod_{1\leq j<k\leq{K-1}}\zeta(1+w_{j}+w_{k})\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}}{\prod_{k=1}^{K-1}\zeta(1+w_{k}+\gamma)\prod_{n=1}^{K-1}(w_{n}-\alpha)(w_{n}+\alpha)}\right)$
×$\times A_{E}(w_{1},\ldots,w_{K-1},\pm\alpha;\gamma)\prod_{k=1}^{K-1}g\left(-\frac{w_{k}}{2}\right)\prod_{j=1}^{K-1}\left[\zeta(1+w_{j}\pm\alpha)(\alpha^{2}-w_{j}^{2})^{2}\right]$
×$\qquad\times dw_{1}\cdots dw_{K-1}.$
@@
×$\times\left(\frac{\prod_{1\leq j<k\leq{K-1}}\zeta(1+\tfrac{w_{j}}{\mathcal{N}_{X}}+\tfrac{w_{k}}{\mathcal{N}_{X}})\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}}{\prod_{k=1}^{K-1}\zeta(1+\tfrac{w_{k}}{\mathcal{N}_{X}}+\gamma)}\right)$
×$\times A_{E}(\tfrac{w_{1}}{\mathcal{N}_{X}},\ldots,\tfrac{w_{K-1}}{\mathcal{N}_{X}},\pm\alpha;\gamma)\prod_{k=1}^{K-1}g\left(-\frac{w_{k}}{2\mathcal{N}_{X}}\right)\prod_{j=1}^{K-1}\left[\zeta(1+\tfrac{w_{j}}{\mathcal{N}_{X}}\pm\alpha)(\big{(}\tfrac{w_{j}}{\mathcal{N}_{X}}\big{)}^{2}-\alpha^{2})\right]$
×$\qquad\times dw_{1}\cdots dw_{K-1}.$
@@
×$\times\oint_{0}\cdots\oint_{0}\exp\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}\sum_{k=1}^{K-1}w_{k}\right)\prod_{1\leq j<k\leq K-1}\left(\frac{1}{w_{j}+w_{k}}\right)\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}$
×$\times\frac{\zeta(1\pm\alpha)^{K-1}}{\zeta(1+\gamma)^{K-1}}A_{E}(0,\ldots,0,\pm\alpha;\gamma)\prod_{n=1}^{K-1}w_{n}^{3-2K}dw_{n}(1+\mathcal{O}(\mathcal{N}_{X}^{-1})).$
@@
[…]
@@
×$\qquad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{ccccc}\frac{1}{\Gamma(2K-3)}&\frac{1}{\Gamma(2K-5)}&\frac{1}{\Gamma(2K-7)}&\cdots&\frac{1}{\Gamma(1)}\\ \frac{1}{\Gamma(2K-4)}&\frac{1}{\Gamma(2K-6)}&\frac{1}{\Gamma(2K-8)}&\cdots&\frac{1}{\Gamma(0)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{1}{\Gamma(K-1)}&\frac{1}{\Gamma(K-3)}&\frac{1}{\Gamma(K-5)}&\cdots&\frac{1}{\Gamma(3-K)}\end{array}\right)$ (195)
@@
×$\times\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}\right)^{(K-1)(K-2)/2}\mathcal{M}^{SO(2N)}(-1)^{K-1}\mathcal{N}_{X}^{(K-1)(K-2)/2}(\pi i)$
×$\times\Bigg{(}\frac{\exp(\mathcal{N}_{d}\alpha)\zeta(1+2\gamma)}{\zeta(1+\gamma)^{K-1}\zeta(1+\alpha+\gamma)}\zeta(1+\alpha)^{K-1}A_{E}(0,\ldots,0,\alpha;\gamma)$
@@
×$\times\Bigg{(}\frac{\zeta(1+2\gamma)}{\zeta(1+\gamma)^{K-1}\zeta(1+\alpha+\gamma)}\zeta(1+\alpha)^{K-1}A_{E}(0,\ldots,0,\alpha;\gamma)$
@@
×$\times\frac{G(r+1)\sqrt{\Gamma(2r+1)}}{\sqrt{G(2r+1)\Gamma(r+1)}}\left(1+\mathcal{O}\left(\mathcal{N}_{X}^{-1}\right)\right).$
@@
$\mathcal{U}_{E}(\mathcal{N}_{d},r,\phi)$ ::=:= $\tilde{A}_{E}^{1}(\mathrm{i}\phi)+\tilde{A}_{E}(\mathrm{i}\phi,\mathrm{i}\phi)\frac{r\zeta^{\prime}(1+\mathrm{i}\phi)}{\zeta(1+\mathrm{i}\phi)}-\tilde{A}_{E}(\mathrm{i}\phi,\mathrm{i}\phi)\frac{\zeta^{\prime}(1+2\mathrm{i}\phi)}{\zeta(1+2\mathrm{i}\phi)}$
@@ 7 One level density for quadratic twists of elliptic curve $L$-functions
For the $(1-c)$ integral, we change variables $s\to 1-s$ so that we can write it as an integral over $(c)$:
@@
We then replace the ratio of $L$-functions with the result in (6), to find that for the integral over $(1-c)$, we get:
@@
×$\left.\times\mathcal{V}(\mathcal{N}_{d},r)\frac{\mathrm{e}^{-\xi r}}{r}f(\phi)\mathrm{d}r\right]\mathrm{d}\phi.$ (213)
@@
×$\left.\times\mathcal{V}_{E}(\mathcal{N}_{d},r)\frac{\mathrm{e}^{-\xi r}}{r}f(\phi)\right]_{r=j}\mathrm{d}\phi.$ (214)

### v2

@@ 6 A mixed moment of elliptic curve $L$-functions
In [14], Conrey, Farmer and Zirnbauer derived the following ratios conjecture. It is not stated explicitly in that paper, but is mentioned after their equation (6.31) as following from their Conjecture 5.3 and their Lemma 6.8 in analogy with (6.31). There are possibly some typos in that paper, so another resource is [35] where the conjecture is stated at Conjecture 5.1 using a slightly different set notation and giving some potential conditions on $\alpha$ and $\gamma$ including ${\rm Re}{\gamma}\gg\frac{1}{\log X}$.
@@
$\times\oint\left(\frac{M|d|^{2}}{4\pi^{2}}\right)^{\tfrac{1}{2}\sum_{k=1}^{K}\left(w_{k}\right)}\left(\frac{H_{\alpha,\gamma}(w_{1},\dots,w_{K})\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\displaystyle\prod_{k=1}^{K}w_{k}}{\displaystyle\prod_{j=1}^{K}\prod_{k=1}^{K}(w_{j}-\alpha_{k})(w_{j}+\alpha_{k})}\right)\mathrm{d}w_{1}\dots\mathrm{d}w_{K},$ (163)
@@
$Y_{E}(\alpha;\gamma):=\$$\frac{\displaystyle\prod_{j<k\leq K}\zeta(1+\alpha_{j}+\alpha_{k})\zeta(1+2\gamma)}{\displaystyle\prod_{k=1}^{K}\zeta(1+\alpha_{k}+\gamma)}.$ (165)
@@
$A_{E}(\alpha,\gamma)$ $:=$ $\frac{\displaystyle\prod_{1\leq j<k\leq K}(1-11^{-1-\alpha_{j}-\alpha_{k}})(1-11^{-1-2\gamma})}{\displaystyle\prod_{k=1}^{K}(1-11^{-1-\alpha_{k}-\gamma})}$ (168)
$\times\left(\frac{(1-\lambda(11)11^{-1/2-\gamma})}{\displaystyle\prod_{k=1}^{K}(1-\lambda(11)11^{-1/2-\alpha_{k}})}\right)$
$\times\displaystyle\prod_{p\neq 11}\left[\frac{\displaystyle\prod_{1\leq j<k\leq K}(1-p^{-1-\alpha_{j}-\alpha_{k}})(1-p^{-1-2\gamma})}{\displaystyle\prod_{k=1}^{K}(1-p^{-1-\alpha_{k}-\gamma})}\right$.
[…]
@@
$\times\oint\exp\left(\mathcal{N}_{d}\sum_{k=1}^{K}w_{k}\right)\left(\frac{H_{\alpha,\gamma}(w_{1},\dots,w_{K})\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\displaystyle\prod_{n=1}^{K}w_{n}^{3-2K}}{\prod_{n=1}^{K}(w_{n}-\alpha)(w_{n}+\alpha)}\right)\mathrm{d}w_{1}\dots\mathrm{d}w_{K}.$
This has exactly the same polar structure as (2) - it behaves like the equivalent integral for $SO(2N)$. In the above we have defined $\mathcal{N}_{d}=\log\left(\frac{\sqrt{M}d}{2\pi}\right)$ to make the comparison easier. The $\mathcal{N}$ associated with the largest $d$ in the sum, $\mathcal{N}_{X}$, can be thought to be equivalent to the matrix size $N$ in the random matrix version. Note that in the random matrix case we found both the leading order term and the next to leading order term in $N$, while here, for simplicity, we shall just work with the leading order term in $\mathcal{N}_{X}$. Acting in analogy to the random matrix case at (2) we know that we can shrink the contours of integration onto the poles, with narrow necks connecting them. As we showed in Section 2.1, the integral will cancel on either side of the necks, leaving a sum over three circular contours for each $w$ variable. The only terms which contribute to this sum will have the residue from one $w_{i}$ evaluated at $\pm\alpha$, and the rest evaluated at 0. Using this fact, and in analogy with (2.1.3), we get
@@
$\times\left(\frac{\prod_{1\leq j<k\leq{K-1}}\zeta(1+w_{j}+w_{k})\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}}{\prod_{k=1}^{K-1}\zeta(1+w_{k}+\gamma)\prod_{n=1}^{K-1}(w_{n}-\alpha)(w_{n}+\alpha)}\right)\prod_{k=1}^{K-1}g\left(-\frac{w_{k}}{2}\right)$
$\times\oint_{\pm\alpha}A_{E}(w_{1},\ldots,w_{K};\gamma)g\left(\frac{\alpha-w_{K}}{2}\right)\prod_{j=1}^{K-1}\left[\zeta(1+w_{j}+w_{K})(w_{K}^{2}-w_{j}^{2})^{2}\right]$
$\qquad\times\frac{\exp(\mathcal{N}_{d}w_{K})w_{K}^{3-2K}}{\zeta(1+w_{K}+\gamma)(w_{K}-\alpha)(w_{K}+\alpha)}dw_{K}\;dw_{1}\cdots dw_{K-1}.$
@@
$\times\left(\frac{\prod_{1\leq j<k\leq{K-1}}\zeta(1+w_{j}+w_{k})\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}}{\prod_{k=1}^{K-1}\zeta(1+w_{k}+\gamma)\prod_{n=1}^{K-1}(w_{n}-\alpha)(w_{n}+\alpha)}\right)$
$\times A_{E}(w_{1},\ldots,w_{K-1},\pm\alpha;\gamma)\prod_{k=1}^{K-1}g\left(-\frac{w_{k}}{2}\right)\prod_{j=1}^{K-1}\left[\zeta(1+w_{j}\pm\alpha)(\alpha^{2}-w_{j}^{2})^{2}\right]$
$\qquad\times dw_{1}\cdots dw_{K-1}.$
@@
$\times\left(\frac{\prod_{1\leq j<k\leq{K-1}}\zeta(1+\tfrac{w_{j}}{\mathcal{N}_{X}}+\tfrac{w_{k}}{\mathcal{N}_{X}})\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}}{\prod_{k=1}^{K-1}\zeta(1+\tfrac{w_{k}}{\mathcal{N}_{X}}+\gamma)}\right)$
$\times A_{E}(\tfrac{w_{1}}{\mathcal{N}_{X}},\ldots,\tfrac{w_{K-1}}{\mathcal{N}_{X}},\pm\alpha;\gamma)\prod_{k=1}^{K-1}g\left(-\frac{w_{k}}{2\mathcal{N}_{X}}\right)\prod_{j=1}^{K-1}\left[\zeta(1+\tfrac{w_{j}}{\mathcal{N}_{X}}\pm\alpha)(\big{(}\tfrac{w_{j}}{\mathcal{N}_{X}}\big{)}^{2}-\alpha^{2})\right]$
$\qquad\times dw_{1}\cdots dw_{K-1}.$
@@
$\times\oint_{0}\cdots\oint_{0}\exp\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}\sum_{k=1}^{K-1}w_{k}\right)\prod_{1\leq j<k\leq K-1}\left(\frac{1}{w_{j}+w_{k}}\right)\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}$
$\times\frac{\zeta(1\pm\alpha)^{K-1}}{\zeta(1+\gamma)^{K-1}}A_{E}(0,\ldots,0,\pm\alpha;\gamma)\prod_{n=1}^{K-1}w_{n}^{3-2K}dw_{n}(1+\mathcal{O}(\mathcal{N}_{X}^{-1})).$
@@
$(K-1)!\left|\begin{array}[]{cccc}\scriptscriptstyle{\oint w_{1}^{3-2K}\exp\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}w_{1}\right)dw_{1}}&\scriptscriptstyle{\oint w_{1}^{5-2K}\exp\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}w_{1}\right)dw_{1}}&\cdots&\scriptscriptstyle{\oint w_{1}^{-1}\exp\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}w_{1}\right)dw_{1}}\\ \vdots&\vdots&\ddots&\vdots\\ \scriptscriptstyle{\oint w_{K-1}^{1-K}\exp\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}w_{K-1}\right)dw_{K-1}}&\scriptscriptstyle{\oint w_{K-1}^{3-K}\exp\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}w_{K-1}\right)dw_{K-1}}&\cdots&\scriptscriptstyle{\oint w_{K-1}^{K-3}\exp\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}w_{K-1}\right)dw_{K-1}}\end{array}\right|$ (185)
$=(K-1)!\left|\begin{array}[]{cccc}\scriptscriptstyle{\left(\tfrac{\mathcal{N}_{X}}{\mathcal{N}_{d}}\right)^{4-2K}\oint w_{1}^{3-2K}\exp\left(w_{1}\right)dw_{1}}&\cdots&\scriptscriptstyle{\oint w_{1}^{-1}\exp\left(w_{1}\right)dw_{1}}\\ \vdots&\ddots&\vdots\\ \scriptscriptstyle{\left(\tfrac{\mathcal{N}_{X}}{\mathcal{N}_{d}}\right)^{2-K}\oint w_{K-1}^{1-K}\exp\left(w_{K-1}\right)dw_{K-1}}&\cdots&\scriptscriptstyle{\left(\tfrac{\mathcal{N}_{X}}{\mathcal{N}_{d}}\right)^{K-2}\oint w_{K-1}^{K-3}\exp\left(w_{K-1}\right)dw_{K-1}}\end{array}\right|$ (189)
@@
$\qquad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{ccccc}\frac{1}{\Gamma(2K-3)}&\frac{1}{\Gamma(2K-5)}&\frac{1}{\Gamma(2K-7)}&\cdots&\frac{1}{\Gamma(1)}\\ \frac{1}{\Gamma(2K-4)}&\frac{1}{\Gamma(2K-6)}&\frac{1}{\Gamma(2K-8)}&\cdots&\frac{1}{\Gamma(0)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{1}{\Gamma(K-1)}&\frac{1}{\Gamma(K-3)}&\frac{1}{\Gamma(K-5)}&\cdots&\frac{1}{\Gamma(3-K)}\end{array}\right)$ (195)
@@
$\times\left(\frac{\mathcal{N}_{d}}{\mathcal{N}_{X}}\right)^{(K-1)(K-2)/2}\mathcal{M}^{SO(2N)}(-1)^{K-1}\mathcal{N}_{X}^{(K-1)(K-2)/2}(\pi i)$
$\times\Bigg{(}\frac{\exp(\mathcal{N}_{d}\alpha)\zeta(1+2\gamma)}{\zeta(1+\gamma)^{K-1}\zeta(1+\alpha+\gamma)}\zeta(1+\alpha)^{K-1}A_{E}(0,\ldots,0,\alpha;\gamma)$
@@
$\times\Bigg{(}\frac{\zeta(1+2\gamma)}{\zeta(1+\gamma)^{K-1}\zeta(1+\alpha+\gamma)}\zeta(1+\alpha)^{K-1}A_{E}(0,\ldots,0,\alpha;\gamma)$
@@
$\times\frac{G(r+1)\sqrt{\Gamma(2r+1)}}{\sqrt{G(2r+1)\Gamma(r+1)}}\left(1+\mathcal{O}\left(\mathcal{N}_{X}^{-1}\right)\right).$
@@
$\mathcal{U}_{E}(\mathcal{N}_{d},r,\phi)$ $:=$ $\tilde{A}_{E}^{1}(\mathrm{i}\phi)+\tilde{A}_{E}(\mathrm{i}\phi,\mathrm{i}\phi)\frac{r\zeta^{\prime}(1+\mathrm{i}\phi)}{\zeta(1+\mathrm{i}\phi)}-\tilde{A}_{E}(\mathrm{i}\phi,\mathrm{i}\phi)\frac{\zeta^{\prime}(1+2\mathrm{i}\phi)}{\zeta(1+2\mathrm{i}\phi)}$
@@ 7 One level density for quadratic twists of elliptic curve $L$-functions
For the (1c)(1-c) integral, we change variables $s\to 1-s$ so that we can write it as an integral over $(c)$:
@@
We then replace the ratio of $L$-functions with the result in (6), to find that for the integral over (1c)(1-c), we get:
@@
$\left.\times\mathcal{V}(\mathcal{N}_{d},r)\frac{\mathrm{e}^{-\xi r}}{r}f(\phi)\mathrm{d}r\right]\mathrm{d}\phi.$ (213)
@@
$\left.\times\mathcal{V}_{E}(\mathcal{N}_{d},r)\frac{\mathrm{e}^{-\xi r}}{r}f(\phi)\right]_{r=j}\mathrm{d}\phi.$ (214)